# LLMs for Literature Review: Are we there yet?

### v1

@@ Abstract
Literature reviews are an essential component of scientific research, but they remain time-intensive and challenging to write, especially due to the recent influx of research papers. This paper explores the zero-shot abilities of recent Large Language Models (LLMs) in assisting with the writing of literature reviews based on an abstract. We decompose the task into two components: (1) Retrieving related works given a query abstract and (2) Writing a literature review based on the retrieved results. We analyze how effective LLMs are for both components. For retrieval, we introduce a novel two-step search strategy that first uses an LLM to extract meaningful keywords from the abstract of a paper and then retrieves potentially relevant papers by querying an external knowledge base. Additionally, we study a prompting-based re-ranking mechanism with attribution and show that re-ranking doubles the normalized recall compared to naive search methods, while providing insights into the LLM’s decision-making process. In the generation phase, we propose a two-step approach that first outlines a plan for the review and then executes steps in the plan to generate the actual review. To evaluate different LLM-based literature review methods, we create test sets from arXiv papers using a protocol designed for rolling use with newly released LLMs to avoid test set contamination in zero-shot evaluations. We release this evaluation protocol to promote additional research and development in this regard. Our empirical results suggest that LLMs show promising potential for writing literature reviews when the task is decomposed into smaller components of retrieval and planning. Particularly, we find that combining keyword-based and document-embedding-based search improves precision and recall during retrieval by 10% and 30% respectively, compared to using either of the methods in isolation. Further, we demonstrate that our planning-based approach achieves higher-quality reviews by minimizing hallucinated references in the generated review by 18-26% compared to existing simpler LLM-based generation methods. Code at github.com.
@@ 1 Introduction
Writing a literature review—finding, citing, and contextualizing relevant prior work—is a fundamental requirement of scientific writing. It is a complex task and can be broken down into two broad sub-tasks: 1) Finding relevant papers and 2) Generating a related work section that situates the proposed research within the existing literature. This challenge is further amplified in the field of machine learning, with thousands of relevant papers appearing every month on ArXiv alone. [^1] Over 4,000 ML papers were submitted to ArXiv in October 2024: https://arxiv.org/list/cs.LG/2024-10 We explore the utility and potential of large language models (LLMs), in combination with retrieval mechanisms, to assist in generating comprehensive literature reviews for scientific papers.
@@
Specifically, we investigate the use of LLMs to generate the related work section of a paper based on its abstract. We use the term abstract loosely, not necessarily to refer to the actual abstract of the paper, but rather a textual passage that captures a concise summary of the paper’s key contributions and scope. Using the abstract as input allows our system to target the central ideas of the paper without requiring the complete manuscript, which is often continuously evolving in the early stages of writing. While our experiments focus on using the abstract, our framework is designed to be flexible and can make use of the entire manuscript as it evolves, albeit at a larger computational cost and the need to use models that support long context windows. This approach provides valuable early-stage insights for authors seeking preliminary references to shape their work, with the capacity to seamlessly incorporate additional information as the manuscript develops.
@@
The architecture of our framework is illustrated in Figure 1. Correspondingly, in this work: 1) We introduce an LLM-based approach to retrieve relevant papers, where we first extract the keywords from an abstract or research idea paragraph using an LLM, and then feed these keywords to a keyword based search tool – we experiment with Google search and Semantic Scholar. We optionally also transform the abstract or idea into an embedding and use an embedding based search procedure as well. 2) We then employ a prompting-based approach to rank the set of retrieved candidate papers based on their relevance to the query abstract, also requiring the LLM to attribute the relevance to specific excerpts in the candidate papers. We explore multiple re-ranking and aggregation strategies 3) To generate the literature review, we select top-$k$ papers from the ranked list and prompt the LLM to generate the related work section based on the query abstract and the abstracts of the selected papers. 4) Additionally, we examine the effectiveness of providing a writing plan to the LLM that specifies which papers to cite at various points in the literature review. These plans can be generated entirely by the LLM, by the user, or a combination of the two. These plans serve as an intermediate representation giving the user more control over the organizational structure of the literature review.
@@
Figure 1: A schematic diagram of our framework, where: 1) Relevant prior work is retrieved using keyword and embedding based search. 2) LLMs re-rank results to find the most relevant prior work. 3) Based on these papers and the user abstract or idea summary, an LLM generates a literature review, (4) optionally controlled by a sentence plan.
@@
Our most powerful model involves multiple innovations, where we use LLMs in multiple ways, namely for generating search queries, re-ranking search results, and attribution. We summarize the main contributions of our work as follows:
@@
  • To answer the key question that our paper poses, we present a data collection protocol and multiple instances of using it to collect arXiv papers. Critically, our protocol is based on using the most recent past month of arXiv papers in a rolling manner with the goal of avoiding test-set contamination when evaluating the most recent LLMs for literature review-related tasks. We then use this protocol to perform extensive retrieval and literature review generation experiments. We release both our datasets and our code to the community.
@@
  • We propose a novel LLM-based pipeline for the task of interactive literature review writing, which we decompose into two distinct components: retrieval and generation. This also facilitates more controlled studies investigating alternative LLM-based approaches for these sub-tasks. Our experiments focus on evaluating fully automated variants of these sub-tasks, but our framing of the problem and our proposed solutions are easily integrated into scenarios where human users interact with systems that assist them. To the best of our knowledge our decomposition of the problem into sub-tasks along with our proposed solutions to them and our framing of this assistive scenario is novel.
@@
  • We make multiple contributions that improve the retrieval phase of the two step process. Firstly, we propose and evaluate a novel strategy that first uses an LLM to extract meaningful keywords from the abstract of a paper and then queries different types of external sources to retrieve potentially relevant papers. Second, we compare and combine the aforementioned LLM generated keyword search techniques, with document embedding based retrieval methods. Third, we propose and examine a wide variety of search re-ranking techniques. Among these we propose a novel prompting for attribution approach, which we find to empirically improve the relevance of retrieved literature, while also improving reliability, providing insights and improved transparency into the decision-making process of LLMs when used to rerank results. Going even further we examine debate prompted LLMs for aggregating and re-ranking keyword search and embedding search results. Our experiments show that combining these ideas improves precision and normalized recall by 10% and 30%, respectively, compared to standard retrieval methods.
@@
  • For text generation, we propose and examine a plan-based retrieval augmented approach to writing of literature reviews. By using a plan and conditioning on retrieved context we provide a user greater control over generated content and our experiments show that this approach can improve the quality of the generated literature reviews substantially. We evaluate the approach using both automated metrics and human assessments and show that our method generates higher-quality reviews as measured by ROUGE scores and as asessed by human evaluation. Our approach also reduces hallucinations by 18-26%.
@@ 2 Related Work
The concept of literature review generation using large language models (LLMs) is built upon the foundation laid by the Multi-XScience dataset proposed by Lu et al. (2020). This dataset paves the way for the challenging task of multi-document summarization, specifically focusing on generating the related-work section of a scientific paper. As underlined by Lu et al. (2020), this approach favors abstractive models, which are well suited for the task. However, unlike the approach suggested by Lu et al. (2020), our work introduces the use of intermediate plans to improve the quality of generated literature reviews. The empirical evidence presented in our study shows that our novel strategy outperforms the vanilla zero-shot generation previously championed by the Multi-XScience dataset (Lu et al., 2020). (Note: This paragraph was entirely generated by GPT-4 following plan-based generation. [^2] We use the plan: Please generate 5 sentences in 60 words. Cite @cite_1 at line 1, 3 and 5. We postprocess to replace delexicalized tokens with latex commands. Outputs from other models are compared later in Appendix (Tables 12 and 13).)
@@
Recent efforts have explored the application of proprietary and open-source LLMs for ranking  (Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023a; b), where the LLM is passed a combined list of passages directly as input and prompted to rank them based on a criteria. Notably, only top-k candidates are passed as input to the LLM for re-ranking (Zhang et al., 2023). In our work, we instruct an LLM to output an ordering of the different candidate papers (e.g. $[3]>[8]>[6]$) in descending order based on the relevance to the user-provided abstract.
@@ 3 Retrieval of Related Work
3.1 Dataset Construction.
We create two datasets that contain papers posted on ArXiv in August and December 2023 respectively, starting with 1,000 papers from each month. We use ArXiv wrapper in Python [^4] https://pypi.org/project/arxiv/ to create RollingEval datasets. We then filter out papers for which we were able to retrieve 100 relevant paper results using LLM summarized keywords. We query the Semantic Scholar API available through the Semantic Scholar Open Data Platform (Lo et al., 2020; Kinney et al., 2023) to search for the relevant papers. To get google search results, we use SERP API [^5] https://serpapi.com/, specifically conditioned to leverage the "site:arxiv.org" parameter. This approach ensures the retrieval of search results are sourced solely from arXiv.org.
To combine results from multiple queries, we take the equal number of top results from each query to get a total of 100 papers. We took caution to avoid duplicate results from different queries. In case we are not able to retrieve sufficient number of results from a query, we then take the equal number from the rest of the queries. This way we ensure that we always retrieve a candidate pool of 100 possible related work for each query paper. We pass these papers as queries to our literature review generation pipeline.
@@
Semantic Scholar API (Multiple queries) + SPECTER2 9.8% 8.2%
Table 1: We created two datasets to measure the efficacy of search using LLM generated keywords - RollingEval-Aug and RollingEval-Dec. We evaluate the % of the ground truth references covered in the top 100 search list using LLM based keyword search and different academic search engines. We note that using multiple queries gives us an edge over using a single query; and we obtain a similar coverage for Semantic Scholar and the SERP API based Google Search.
We now describe our two-step retrieval mechanism also demonstrated in Algorithm 1.
@@ 3.2 Retrieving Candidate Papers
To retrieve related work for a given paper abstract, first, for each query abstract in the dataset, we prompt an LLM to generate keywords that we use as queries for a search API (refer to Figure 14 in the Appendix for the detailed prompt used for this task). Importantly, we add a timestamp filter to the search API to retrieve papers published stricly before the publication date of the query paper. In addition to evaluating multiple search engines, we also experiment with generating multiple queries  [^6] In our experiments, we generate three queries for each abstract. from an LLM and various heuristics to combine the search results from each query (see Appendix D). We evaluate multiple general and academic search engines on the quality of the retrieved papers using coverage, which we define as the percentage of ground-truth papers retrieved by the search engine. Table 1 shows the coverage for different search engines and query heuristics. We note that using multiple queries achieves the highest coverage with comparable results for Semantic Scholar search and SERP API. However, at best, we retrieve just under 7% of the ground truth papers. The low retrieval percentage of just under 7% can be attributed to several factors. First, the task of finding related work for a given paper is inherently challenging due to the diverse styles and methods authors use in literature reviews. This stylistic variability means that a one-size-fits-all approach, such as generating search keywords and using a search engine, might not capture the nuanced criteria that a human expert would apply. Additionally, our retrieval process operates in a constrained setting, generating search keywords based solely on the abstract of the paper. In theory, including more of the paper—such as the introduction or methodology—could provide richer context and lead to higher retrieval rates. However, this would not align with our intended use case: supporting researchers in the early stages of drafting when only limited, unpolished material is available.
We also explore an embedding-based strategy for retrieval using SPECTER embeddings (Cohan et al., 2020). SPECTER uses a contrastive learning approach to train a SciBERT (Beltagy et al., 2019) model based on a triplet loss to discriminate between related versus unrelated papers. SciBERT is a BERT-like transformer model (Devlin et al., 2018) trained on scientific text. More recently, Singh et al. (2022) further extended the contrastive learning approach of SPECTER to better deal with multiple tasks of relevance to scientific papers, including citation prediction, leading to SPECTER2. Since SPECTER2 provides a large knowledge base of 150M document embeddings of scientific articles, we explore embedding-based retrieval of candidate papers, where we obtain the top-$k$ papers based on the cosine similarity between the query abstract and the corpus of candidates papers in the SPECTER2 dataset.

### v2

# LitLLMs, LLMs for Literature Review: Are we there yet?

@@ Abstract
Literature reviews are an essential component of scientific research, but they remain time-intensive and challenging to write, especially due to the recent influx of research papers. This paper explores the zero-shot abilities of recent Large Language Models (LLMs) in assisting with the writing of literature reviews based on an abstract. We decompose the task into two components: (1) Retrieving related works given a query abstract and (2) Writing a literature review based on the retrieved results. We analyze how effective LLMs are for both components. For retrieval, we introduce a novel two-step search strategy that first uses an LLM to extract meaningful keywords from the abstract of a paper and then retrieves potentially relevant papers by querying an external knowledge base. Additionally, we study a prompting-based re-ranking mechanism with attribution and show that re-ranking doubles the normalized recall compared to naive search methods while providing insights into the LLM’s decision-making process. In the generation phase, we propose a two-step approach that first outlines a plan for the review and then executes steps in the plan to generate the actual review. To evaluate different LLM-based literature review methods, we create test sets from arXiv papers using a protocol designed for rolling use with newly released LLMs to avoid test set contamination in zero-shot evaluations. We release this evaluation protocol to promote additional research and development in this regard. Our empirical results suggest that LLMs show promising potential for writing literature reviews when the task is decomposed into smaller components of retrieval and planning. Particularly, we find that combining keyword-based and document-embedding-based search improves precision and recall during retrieval by 10% and 30%, respectively, compared to using either of the methods in isolation. Further, we demonstrate that our planning-based approach achieves higher-quality reviews by minimizing hallucinated references in the generated review by 18-26% compared to existing simpler LLM-based generation methods. Our project page including a demonstration system and toolkit can be accessed here: https://litllm.github.io.
@@ 1 Introduction
Writing a literature review—finding, citing, and contextualizing relevant prior work—is a fundamental scientific writing requirement. When writing manuscripts, scientists must situate their proposed ideas within the existing literature. Writing a good literature review is a complex task which can be broken down into two broad sub-tasks: 1) Finding relevant papers and 2) Generating a related work section to discuss the proposed research given prior works. This challenge is further amplified in fields such as machine learning, where thousands of relevant papers appear every month on arXiv alone. [^1] E.g. over 4,000 ML papers were submitted to arXiv in October 2024: https://arxiv.org/list/cs.LG/2024-10 We explore the utility and potential of large language models (LLMs), in combination with retrieval mechanisms, to assist in generating comprehensive literature reviews for scientific papers.
@@
Specifically, we investigate using LLMs to generate a paper’s related work section based on its abstract. We use the term abstract loosely, not necessarily to refer to the actual abstract of the paper but rather to a textual passage that captures a concise summary of the paper’s key contributions and scope. Using the abstract as input allows our system to target the central ideas of the paper without requiring the complete manuscript, which is often continuously evolving in the early stages of writing. While our experiments focus on using the abstract, our framework is designed to be flexible. It can use the entire manuscript as it evolves, albeit at a higher computational cost and the need to use models that support longer context windows. This approach provides valuable early-stage insights for authors seeking preliminary references to shape their work, with the capacity to seamlessly incorporate additional information as the manuscript develops.
@@
The architecture of our framework is illustrated in Figure 1, where we further decompose each of the two above tasks into two subtasks. In this work: 1) We introduce an LLM-based approach to retrieve relevant papers, where we first extract the keywords from an abstract or research idea paragraph using an LLM and then feed these keywords to a keyword-based search tool — we experiment with Google search and Semantic Scholar. We optionally also transform the abstract or idea into an embedding and use an embedding-based search procedure. 2) We then employ a prompting-based approach to rank the set of retrieved candidate papers based on their relevance to the query abstract, also requiring the LLM to attribute the relevance to specific excerpts in the candidate papers. We explore multiple re-ranking and aggregation strategies. 3) To generate the literature review, we select top-$k$ papers from the ranked list and prompt the LLM to generate the related work section based on the query abstract and the abstracts of the selected papers. 4) Additionally, we examine the effectiveness of providing a writing plan to the LLM that specifies which papers to cite at various points in the literature review. These plans can be generated entirely by the LLM, by the user, or a combination of the two. These plans serve as an intermediate representation giving the user more control over the organizational structure of the literature review.
@@
Figure 1: A schematic diagram of our framework, where: 1) Relevant prior work is retrieved using keyword and embedding-based search. 2) LLMs re-rank results to find the most relevant prior work. 3) Based on these papers and the user abstract or idea summary, an LLM generates a literature review, 4) optionally controlled by a sentence plan.
@@
The complete framework involves multiple innovations, where we use LLMs in multiple ways, namely for generating search queries, re-ranking search results, and attribution. We summarize the main contributions of our work as follows:
@@
  • To answer the key question that our paper poses, we present a data collection protocol and multiple instances of using it to collect arXiv papers. Critically, our protocol is based on using the most recent month of arXiv papers in a rolling manner with the goal of avoiding test-set contamination when evaluating the most recent LLMs for literature review-related tasks. We then use this protocol to perform extensive retrieval and literature review generation experiments. We release both our datasets and our code to the community.
@@
  • We propose a novel LLM-based pipeline for the task of interactive literature review writing, which we decompose into two distinct components: retrieval and generation. This also facilitates more controlled studies investigating alternative LLM-based approaches for these sub-tasks. Our experiments focus on evaluating fully automated variants of these sub-tasks, but our framing of the problem and our proposed solutions are easily integrated into scenarios where human users interact with systems that assist them. To the best of our knowledge, our decomposition of the problem into sub-tasks, along with our proposed solutions to them, and our framing of this assistive scenario is novel.
@@
  • We make multiple contributions that improve the retrieval phase of the two-step process. First, we propose and evaluate a novel strategy that first uses an LLM to extract meaningful keywords from the abstract of a paper and then queries different types of external sources to retrieve potentially relevant papers. Second, we compare and combine the aforementioned LLM-generated keyword search techniques with document embedding-based retrieval methods. Third, we propose and examine a wide variety of search re-ranking techniques. Among these, we propose a novel prompting for attribution approach, which we find to empirically improve the relevance of retrieved literature while also improving reliability, providing insights and improving transparency in the decision-making process of LLMs when used to rerank results. Going even further, we examine debate-prone LLMs for aggregating and re-ranking keyword search and embedding search results. Our experiments show that combining these ideas improves precision and normalized recall by 10% and 30%, respectively, compared to standard retrieval methods.
@@
  • For text generation, we propose and examine a plan-based retrieval augmented approach to writing literature reviews. By using a plan and conditioning on retrieved context, we provide a user greater control over generated content, and our experiments show that this approach can improve the quality of the generated literature reviews substantially. We evaluate the approach using automated metrics and human assessments and show that our method generates higher-quality reviews as measured by ROUGE scores and as assessed by human evaluation. Our approach also reduces hallucinations by 18–26%.
@@ 2 Related Work
We decompose the literature review task into two key sub-tasks: identifying relevant papers and generating the final related work section. This decomposition enhances the likelihood that the LLMs can effectively accomplish the task. We now discuss the relevant literature pertinent to both aspects of the process.
2.1 Ranking and Retrieval
Traditional methods for information retrieval rely on techniques like TF-IDF and BM25 to identify documents that are semantically similar to a given query. More recently, dense vector representations obtained through models like Sentence-BERT (Reimers & Gurevych, 2019) have been shown to improve retrieval accuracy by encoding both query and documents into an embedding space where semantic similarity can be readily computed. The initial retrieval stage often results in a large set of candidate documents, which then need to be re-ranked to obtain an ordered list based on relevance.
Recent efforts have explored the application of proprietary and open-source LLMs for ranking  (Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023a; b), where the LLM is passed a combined list of passages directly as input and prompted to rank them based on a criteria. Notably, only top-k candidates are passed as input to the LLM for re-ranking (Zhang et al., 2023). In our work, we instruct an LLM to output an ordering of the different candidate papers (e.g. $[3]>[8]>[6]$) in descending order based on the relevance to the user-provided abstract. Although existing re-ranking methods improve the ordering of candidate papers, they do not provide explicit justification for the relative rankings assigned. To improve the reliability of the system and offer a clear explanation for the model’s choices, our methodology also incorporates attribution capabilities, allowing us to identify specific textual elements contributing to relevance scores for different candidates. Among works exploring the attribution capabilities in LLMs, Yue et al. (2023) focuses on automatically evaluating whether generated statements are fully supported by cited references. Cohen-Wang et al. (2024) present ContextCite, a method for attributing model generation to context, which can be applied on top of any existing language model to help verify generated statements, improve response quality, and detect poisoning attacks. Gradient-based techniques, such as Integrated Gradients (Sundararajan et al., 2017) and Saliency Maps (Simonyan et al., 2014), measure the contribution of each input token by computing gradients of the output with respect to input features; advances like SmoothGrad (Smilkov et al., 2017) and DeepLIFT (Shrikumar et al., 2017) improve these methods by reducing noise and enhancing accuracy. Perturbation techniques involve modifying or occluding parts of the input and observing changes in the output to infer input significance (Li et al., 2016), utilizing methods like Meaningful Perturbations (Fong & Vedaldi, 2017) and LIME (Ribeiro et al., 2016). Despite advancements, challenges such as attribution leakage (Adebayo et al., 2018), unreliability of saliency methods (Kindermans et al., 2019), and complexities in attributing outputs in large models (Ghorbani et al., 2019) persist. Surveys like Li et al. (2023) discuss current methodologies and inherent challenges, while research by Keeling & Street (2024) examines the theoretical basis for attributing confidence to LLMs, raising concerns about the reliability of experimental assessment techniques. In contrast to the discussed gradient-based attribution methods that are challenging to scale and perturbation-based approaches that need multiple passes through the model, we propose a straightforward prompting-based attribution approach that can be applied to any LLM agent, is readily scalable, and does not require multiple passes through the model.
2.2 Literature Review Generation
@@
The concept of literature review generation using large language models (LLMs) is built upon the foundation laid by the Multi-XScience dataset proposed by Lu et al. (2020). This dataset paves the way for the challenging task of multi-document summarization, specifically focusing on generating the related work section of a scientific paper. As underlined by Lu et al. (2020), this approach favors abstractive models, which are well suited for the task. However, unlike the approach suggested by Lu et al. (2020), our work introduces the use of intermediate plans to improve the quality of generated literature reviews. The empirical evidence presented in our study shows that our novel strategy outperforms the vanilla zero-shot generation previously championed by the Multi-XScience dataset (Lu et al., 2020). (Note: This paragraph was entirely generated by GPT-4 following plan-based generation. [^2] We use the plan: Please generate 5 sentences in 60 words. Cite @cite_1 at line 1, 3 and 5. We postprocess to replace delexicalized tokens with latex commands. Outputs from other models are compared later in Appendix (Tables 13 and 14).)
@@ 3 Retrieval of Related Work
3.1 Dataset Construction
We create two datasets that contain papers posted on arXiv in August and December 2023, respectively, starting with 1,000 papers from each month. We use the arXiv wrapper in Python [^4] https://pypi.org/project/arxiv/ to create RollingEval datasets. We then filter out papers for which we were not able to retrieve 100 relevant paper results using LLM summarized keywords. We query the Semantic Scholar API available through the Semantic Scholar Open Data Platform (Lo et al., 2020; Kinney et al., 2023) to search for the relevant papers. To get Google search results, we use SERP API, [^5] https://serpapi.com/ specifically conditioned to leverage the “site:arxiv.org” parameter. This approach ensures the retrieval of search results are sourced solely from arXiv.org.
To combine results from multiple queries, we take the equal number of top results from each query to get a total of 100 papers. We took caution to avoid duplicate results from different queries. In case we are not able to retrieve a sufficient number of results from a query, we then take an equal number from the rest of the queries. This way we ensure that we always retrieve a candidate pool of 100 possible related work for each query paper. We pass these papers as queries to our literature review generation pipeline.
@@ 3.1 Dataset Construction
Semantic Scholar API (Multiple queries) + SPECTER2 9.80 8.20
Table 1: We created two datasets to measure the efficacy of search using LLM-generated keywords - RollingEval-Aug and RollingEval-Dec. We evaluate the % of the ground truth references covered in the top 100 search list using LLM-based keyword search and different academic search engines. We note that using multiple queries gives us an edge over using a single query, and we obtain a similar coverage for Semantic Scholar and the SERP API-based Google Search.
We now describe our two-step retrieval mechanism and provide its pseudo-code in Algorithm 1.
@@ 3.2 Retrieving Candidate Papers
To retrieve related work for a given paper abstract, first, for each query abstract in the dataset, we prompt an LLM to generate keywords that we use as queries for a search API (refer to Figure 14 in the Appendix for the detailed prompt used for this task). Importantly, we add a timestamp filter to the search API to retrieve papers published strictly before the publication date of the query paper. In addition to evaluating multiple search engines, we also experiment with generating multiple queries [^6] In our experiments, we generate three queries for each abstract. from an LLM and various heuristics to combine the search results from each query (see Appendix D). We evaluate multiple general and academic search engines on the quality of the retrieved papers using coverage, which we define as the percentage of ground-truth papers retrieved by the search engine.
Table 1 shows the coverage for different search engines and query heuristics. We note that using multiple queries achieves the highest coverage with comparable results for Semantic Scholar search and SERP API. However, at best, we retrieve just under 7% of the ground truth papers. The low retrieval percentage of just under 7% can be attributed to several factors. First, the task of finding related work for a given paper is inherently challenging due to the diverse styles and methods authors use in literature reviews. This stylistic variability means that a one-size-fits-all approach, such as generating search keywords and using a search engine, might not capture the nuanced criteria that a human expert would apply. Additionally, our retrieval process operates in a constrained setting, generating search keywords based solely on the paper’s abstract. In theory, including more of the paper—such as the introduction or methodology—could provide richer context and lead to higher retrieval rates. However, this would not align with our intended use case: supporting researchers in the early stages of drafting when only limited, unpolished material is available.
We also explore an embedding-based strategy for retrieval using SPECTER embeddings (Cohan et al., 2020). SPECTER uses a contrastive learning approach to train a SciBERT (Beltagy et al., 2019) model based on a triplet loss to discriminate between related versus unrelated papers. SciBERT is a BERT-like transformer model (Devlin et al., 2018) trained on scientific text. More recently, Singh et al. (2022) further extended the contrastive learning approach of SPECTER to better deal with multiple tasks of relevance to scientific papers, including citation prediction, leading to SPECTER2. Since SPECTER2 provides a large knowledge base of 150M document embeddings of scientific articles, we explore embedding-based retrieval of candidate papers, where we obtain the top-$k$ papers based on the cosine similarity between the query abstract and the corpus of candidate papers in the SPECTER2 dataset.